Based on the discuss records from the Olympics the winning distance can be approximated by d=175+1.75t where d is in feet and t=0 corresponds to the year 1948.estimate the year in which the winning toss should be about 260feet
Answer by ankor@dixie-net.com(22740)

Based on the discus records from the Olympics the winning distance can be approximated by d=175+1.75t where
d is in feet and t=0 corresponds to the year 1948.
estimate the year in which the winning toss should be about 260feet
:
d = 260, therefore
175 + 1.75t = 260
1.75t = 260 - 175
1.75t = 85
t = 85/1.75
t = 48.57 ~ 49
:
1948 + 49 = 1997 the yr of the winning toss; 260 ft